OpenStudy (anonymous):

4 sin^2 x - 4 sin x + 1 = 0 solutions within 2pi and 0

OpenStudy (anonymous):

i think it is a perfect square

OpenStudy (anonymous):

I know how to graph, but I'm having issues solving algebraically

OpenStudy (anonymous):

Yes it is (2sin-1)^2

OpenStudy (anonymous):

\[(2\sin(x)-1)^2=0\]

OpenStudy (anonymous):

I got pi/6

OpenStudy (anonymous):

But I know from graphing that it is pi/6 and 5pi/6

OpenStudy (anonymous):

so \[2sin(x)-1=0\\ 2\sin(x)=1\\ \sin(x)=\frac{1}{2}\]

OpenStudy (anonymous):

arcsin(1/2) = pi/6

OpenStudy (anonymous):

yeah but arcsine only gives you the number on the right side of the unit circle

OpenStudy (anonymous):

Then I know I add pi to it to get the other side

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (anonymous):

Reflect over the y?

OpenStudy (anonymous):

aka subtract from \(\pi\)

OpenStudy (anonymous):

Ok thanks man

OpenStudy (anonymous):

Same y though? Isnt cosine x?

OpenStudy (anonymous):

yw, always glad to help a ruthless dictator

OpenStudy (anonymous):

yeah same y value

OpenStudy (anonymous):

And the x's are my solutions

OpenStudy (anonymous):

it was sine right

OpenStudy (anonymous):

Oh oops

OpenStudy (anonymous):

ok maybe that was not a good way to put it same second coordinate on the unit circle

OpenStudy (anonymous):

Makes a lot of sense! Thanks a lot.